OpenStudy (anonymous):

what is the ratio of 5/12 to 7/9?

OpenStudy (vishweshshrimali5):

See, 5/12 : 7/9 = (5/12)/(7/9) = 5 * 9/12*7 = 45/84 = 45 : 84 This should be your answer

OpenStudy (vishweshshrimali5):

Satisfied with the answer !!! :)

OpenStudy (vishweshshrimali5):

If yes then please give me a medal

OpenStudy (anonymous):

45:84 which simplifies to 15:28, the answer in the answer key. thanks.

OpenStudy (vishweshshrimali5):

Mention no phoebeco